# Predictability-Aware Compression and Decompression Framework for Multichannel Time Series Data

### v1

@@ 3.4 Ablation results
4 Conclusion
This paper proposes a novel predictability-aware compression–decompression framework tailored for multichannel time series prediction. By leveraging a periodicity-guided convolution process with circularly orthogonal key matrices, the framework effectively preserves latent temporal dependencies during compression and mitigates reconstruction errors during decompression. Theoretical analyses verify that the introduced compression–decompression components do not dominate the overall computational complexity, while practical designs such as sparse encoding and lightweight residual modules further alleviate the burden on edge devices. Extensive experiments across diverse datasets and prediction backbones demonstrate that our framework consistently achieves lower runtime and improved prediction accuracy, as evidenced by reduced Cobb–Douglas production indices. Furthermore, ablation results confirm the necessity of jointly optimizing both compression and decompression processes. Overall, this work offers a flexible and efficient solution for compressing multichannel time series data without sacrificing predictive performance, paving the way for more resource-efficient forecasting in edge–cloud environments.
Limitations and future work
@@ 4 Conclusion
Broader impacts
The proposed framework significantly reduces computational resources on edge devices and communication requirements on the local network, enabling deployment in bandwidth- or resource-constrained environments. This has practical societal benefits in both developed and underdeveloped areas. For example, in AI smart-city, fast edge predictions improve flow and reduce emissions. Correspondingly, in rural agriculture, lightweight on-device prediction enables timely decision-making despite outdated hardware and poor connectivity.
Appendix A Technical Appendices and Supplementary Material
Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline. There is no page limit for the technical appendices.
A.1 Justification of Predictability Preservation in Compression
The result of the input vector after the seasonal convolution process is computed as follows. For the single element of the single channel $y_{t}^{i}\in Y^{i}$, we have
$y_{t}^{i}=(k_{0}\dots k_{\tau-1})\odot[(x_{1}^{i}\dots x_{\tau-1}^{i})+(x_{\tau}^{i}\dots x_{2\tau-1}^{i})+\dots+(x_{(n-1)\tau}^{i}\dots x_{n\tau-1}^{i})]$ (17)
where $k$ denotes the cyclically shifted version of the seasonal key.
The entire vector $Y^{i}$ form:
$\mathbf{Y^{i}}=\begin{bmatrix}(k_{0}...k_{-\tau+1})\odot[(x_{1}^{i}\dots x_{\tau-1}^{i})+(x_{\tau}^{i}\dots x_{2\tau-1}^{i})+\dots+(x_{(n-1)\tau}^{i}\dots x_{n\tau-1}^{i})]\\ (k_{1}...k_{-\tau+2})\odot[(x_{1}^{i}\dots x_{\tau-1}^{i})+(x_{\tau}^{i}\dots x_{2\tau-1}^{i})+\dots+(x_{(n-1)\tau}^{i}\dots x_{n\tau-1}^{i})]\\ (k_{2}...k_{-\tau+3})\odot[(x_{1}^{i}\dots x_{\tau-1}^{i})+(x_{\tau}^{i}\dots x_{2\tau-1}^{i})+\dots+(x_{(n-1)\tau}^{i}\dots x_{n\tau-1}^{i})]\\ ...\\ (k_{0}...k_{-\tau+1})\odot\overline{\gamma}[(x_{1}^{i}\dots x_{\tau-1}^{i})+(x_{\tau}^{i}\dots x_{2\tau-1}^{i})+\dots+(x_{(n-1)\tau}^{i}\dots x_{n\tau-1}^{i})]\\ ...\\ \end{bmatrix}$ (18)
To evaluate the effectiveness of this encoding in preserving temporal dependencies, we compute the Pearson correlation over the encoded sequence $Y^{i}$:
$R$ $=\frac{\sum_{t}^{t+\tau}(y_{t}^{i}-\bar{y}^{1})(y_{t+\tau}^{i}-\bar{y}^{2})}{\sqrt{\sum_{t}^{t+\tau}(y_{t}^{i}-\bar{y}^{1})^{2}}\sqrt{\sum_{t}^{t+\tau}(y_{t+\tau}^{i}-\bar{y}^{2})^{2}}}$ (19)
$=\frac{\overline{\gamma}[S_{1}^{2}+...+S_{\tau}^{2}]}{\left|\overline{\gamma}\right|[S_{1}^{2}+...+S_{\tau}^{2}]}$
$=1$
where $\overline{\gamma}>0$, $S_{i}$ is computed in the following formulation,
$S_{t}=[y_{t}^{i}-y_{\text{mean}}^{\tau}]^{2}$ (20)
where $y_{\text{mean}}^{\tau}$ represents the mean of the current period encoded elements. If the seasonal convolution aligns with the latent periodicity, the resulting correlation approaches 1, indicating strong predictability preservation.
A.2 Additional reconstruction errors reduction
Decoding process is shown in following formulation:
$\tilde{x}_{t}^{i}=\sum_{z=0}^{n-1}y_{z}^{i}k_{z-t}$ (21)
Accordingly, the decoding procedure can be expressed in matrix form as:
$\tilde{X}^{i}$ $=\begin{bmatrix}\tilde{x}_{0}^{i}&\tilde{x}_{1}^{i}&...&\tilde{x}_{n-1}^{i}\\ \end{bmatrix}$ (22)
$=K\cdot Y$
$=\begin{bmatrix}k_{0}&k_{1}&...&k_{n-1}\\ k_{1}&k_{2}&...&k_{n}\\ ...&...&...&...\\ k_{n-1}&...&...&k_{2n-1}\\ \end{bmatrix}\cdot\begin{bmatrix}y_{0}^{i}\\ y_{1}^{i}\\ ...\\ y_{n-1}^{i}\\ \end{bmatrix}$
$=\begin{bmatrix}k_{0}&k_{1}&...&k_{n-1}\\ k_{1}&k_{2}&...&k_{n}\\ ...&...&...&...\\ k_{n-1}&...&...&k_{2n-1}\\ \end{bmatrix}\cdot\begin{bmatrix}k_{0}x_{0}+k_{-1}x_{1}+...+k_{-n+1}x_{-n+1}\\ k_{1}x_{0}+k_{0}x_{1}+...+k_{-n+2}x_{-n+1}\\ ...\\ k_{n-1}x_{0}+k_{n-2}x_{1}+...+k_{0}x_{-n+1}\\ \end{bmatrix}$
$=\begin{bmatrix}(\sum k^{2})x_{0}^{i}+\eta&(\sum k^{2})x_{1}^{i}+\eta&...&(\sum k^{2})x_{n-1}^{i}+\eta\end{bmatrix}$
$=(\sum k^{2})X^{i}+\eta$
where $\eta$ denotes accumulated noise from neighboring timestamps. The aggregate reconstruction from all channels is:
$\tilde{X}$ $=\begin{bmatrix}k_{0}&k_{1}&...&k_{n-1}\\ k_{1}&k_{2}&...&k_{n}\\ ...&...&...&...\\ k_{n-1}&...&...&k_{2n-1}\\ \end{bmatrix}\cdot\begin{bmatrix}y_{0}^{0}\\ y_{1}^{0}\\ ...\\ y_{n-1}^{0}\\ \end{bmatrix}+...+\begin{bmatrix}k_{0}&k_{1}&...&k_{n-1}\\ k_{1}&k_{2}&...&k_{n}\\ ...&...&...&...\\ k_{n-1}&...&...&k_{2n-1}\\ \end{bmatrix}\cdot\begin{bmatrix}y_{0}^{C-1}\\ y_{1}^{C-1}\\ ...\\ y_{n-1}^{C-1}\\ \end{bmatrix}$ (23)
$=\begin{bmatrix}(\sum k^{2})x_{0}^{0}+(\sum k^{2})x_{0}^{1}+...+(\sum k^{2})x_{0}^{C-1}+C\eta\\ (\sum k^{2})x_{1}^{0}+(\sum k^{2})x_{1}^{1}+...+(\sum k^{2})x_{1}^{C-1}+C\eta\\ ...\\ (\sum k^{2})x_{n-1}^{0}+(\sum k^{2})x_{n-1}^{1}+...+(\sum k^{2})x_{n-1}^{C-1}+C\eta\end{bmatrix}$
In classical HRR, it is assumed that the input vectors the input vector $X$ and keys $K$ follows $N(0,\frac{1}{N})$, allowing $\eta$ to vanish in expectation. However, this assumption does not hold for real-world time series. Moreover, due to the multichannel nature of the problem, recovering a specific channel $X^{i}$ from (23) becomes ill-posed, since cross-channel interference terms $X^{\neq i}$ are also present
To better characterize the noise, we distinguish two sources: Intra-channel interference $\eta$: noise from adjacent timestamps within the same channel; Inter-channel interference: unwanted contributions from other channels at the same timestamp.
To mitigate intra-channel interference, we adopt an Orthogonal Circular Matrix for keys $K$. The characteristic of a circular matrix [5] is that each column is a cyclic shift of the previous column, meaning it is entirely determined by its first row. An Orthogonal Circular Matrix is a special type of circular matrix that satisfies the orthogonality condition $K^{T}K=I$, where $I$ is the identity matrix. (Obtained from two times FFT [24]). The updated aggregated reconstruction becomes:
$\tilde{X}$ $=\begin{bmatrix}(\sum k^{2})x_{0}^{0}+(\sum k^{2})x_{0}^{1}+...+(\sum k^{2})x_{0}^{C-1}+C\eta\\ (\sum k^{2})x_{1}^{0}+(\sum k^{2})x_{1}^{1}+...+(\sum k^{2})x_{1}^{C-1}+C\eta\\ ...\\ (\sum k^{2})x_{n-1}^{0}+(\sum k^{2})x_{n-1}^{1}+...+(\sum k^{2})x_{n-1}^{C-1}+C\eta\end{bmatrix}$ (24)
$=(\sum k^{2})\cdot\begin{bmatrix}x_{0}^{0}+x_{0}^{1}+...+x_{0}^{C-1}\\ x_{1}^{0}+x_{1}^{1}+...+x_{1}^{C-1}\\ ...\\ x_{n-1}^{0}+x_{n-1}^{1}+...+x_{n-1}^{C-1}\end{bmatrix}$
A.3 Information bottleneck
$\mathcal{L}_{\text{IB}}=I(T;X)-\beta I(T;Y)$ (25)
For $I(T;X)$
$Y=KX$ (26)
$X^{(i)}=s^{(i)}+\varepsilon^{(i)},\quad\varepsilon^{(i)}\sim\mathcal{N}(0,\sigma^{2}I),$ (27)
$Y=\underbrace{\sum_{i=1}^{C}Ks^{(i)}}_{\text{signal}}+\underbrace{\sum_{i=1}^{C}K\varepsilon^{(i)}}_{\text{Gaussian noise}}.$ (28)
$I(T;X)\leq\frac{1}{2}\log\left(1+\frac{\sigma_{s}^{2}}{\sigma_{n}^{2}}\right)=\frac{1}{2}\log(1+\mathrm{SNR})$ (29)
For $I(T;X)$
$\log p(y\mid t)=\log q(y\mid t)+\log\frac{p(y\mid t)}{q(y\mid t)},$ (30)
$I(T;Y)=\mathbb{E}_{p(t,y)}\left[\log q(y\mid t)\right]+\underbrace{\mathbb{E}_{p(t,y)}\left[\log\frac{p(y\mid t)}{q(y\mid t)}\right]}_{\text{KL divergence}\geq 0}+H(Y).$ (31)
$I(T;Y)\geq\mathbb{E}_{p(t,y)}\left[\log q(y\mid t)\right]+H(Y)\geq-H(\hat{Y})+H(Y)=0$ (32)
Therefore,
$\mathcal{L}_{\text{IB}}\leq\frac{1}{2}\log(1+\mathrm{SNR})+\beta H(\hat{Y})-\beta H(Y)$ (33)
Based on the normalization applied to the compressed representation, we assume the signal variance $\sigma^{2}_{X}=1$. The term $\beta H(\hat{Y})-\beta H(Y)$ is expected to approach zero during training. Assuming the total noise arises from all channels over a sequence of length $L$, and each noise source follows a distribution $\mathcal{N}(0,\sigma^{2})$, the total noise variance scales with both the number of channels $C$ and the sequence length $L$. Thus, the denominator is proportional to $CL\sigma^{2}$. The upper bound of $\mathcal{L}_{\text{IB}}$ becomes:
$\mathcal{L}_{\text{IB}}\leq\frac{1}{2}\log\left(1+\frac{1}{CL\cdot\sigma^{2}}\right)$ (34)
A.4 Normalization
Followed by Dish-TS [7], a normalization process is applied after compression, and a denormalization is applied before decompression. The normalization process is formulated as:
$Y=[\text{Normal}(Y_{0:\tau}),\text{Normal}(Y_{\tau:2\tau}),...,\text{Normal}(Y_{(\lfloor\frac{N}{\tau}\rfloor-1)\tau:\lfloor\frac{N}{\tau}\rfloor\tau})]$ (35)
Here, $\text{Normal}(\cdot)$ denotes standard normalization, i.e., subtracting the mean and dividing by the standard deviation, following Dish-TS []. This normalization helps stabilize the outputs from different compression periods $\tau$.
The corresponding denormalization step during decompression is defined as:
$\hat{Y}=\text{DeNormal}(\text{Predict}(Y))$ (36)
where $\text{Predict}(\cdot)$ is single channel predictor, $\text{DeNormal}(\cdot)$ is the inverse of the normalization operation.
A.5 Whole processes
Input: Multichannel time series $X\in\mathbb{R}^{C\times N}$, period $\tau$, learning rate $\eta$, total epochs $T$
Output: Final prediction $\hat{X}$, final loss $\mathcal{L}$
Training Phase:
for $\text{epoch}\leftarrow 1$ to $T$ do
       Compression:
       for $i\leftarrow 1$ to $C$ do
             Split $X^{i}$ into segments $X^{i}_{0:\tau},X^{i}_{\tau:2\tau},\dots$
             for each segment do
                   $Y^{i}\leftarrow\text{Encode}(X^{i}_{j\tau:(j+1)\tau},K)$
      $Y\leftarrow\sum_{i=1}^{C}Y^{i}$
      Normalization:
       for each segment $Y_{j\tau:(j+1)\tau}$ do
             $Y_{j\tau:(j+1)\tau}\leftarrow\text{Normal}(Y_{j\tau:(j+1)\tau})$
      Prediction and Denormalization:
       $\hat{Y}\leftarrow\text{DeNormal}(\text{Predict}(Y))$
      Decompression:
       $\hat{X}\leftarrow(K\cdot\hat{Y}/\sum k^{2})+\text{Residual}(K\cdot\hat{Y}/\sum k^{2})$
      Loss Computation:
       $\mathcal{L}\leftarrow\|X-\hat{X}\|^{2}+\alpha(\sum|\text{Residual}|-\sum\hat{X})^{2}+\beta\|\varphi_{Y}-\varphi_{\hat{Y}}\|^{2}$
      Parameter Update:
       $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}$
Testing Phase:
Repeat lines 2–17 without the parameter update step.
Output final $\hat{X}$ and $\mathcal{L}$.
Algorithm 1 Multichannel compression and decompression for time series prediction
To conclude, the overall training and inference workflow is delineated in Algorithm 1 and comprises the following three key stages:
Sparse Temporal Compression and Normalization: Each channel of the input time series is partitioned into non-overlapping segments and encoded via a parameterized compression module. To ensure stability across varying periodicities, segment-wise normalization is subsequently applied.
Unified Prediction and Lightweight Decompression: A shared predictor operates on the aggregated compressed representation. The predicted output is normalized by the energy of the seasonal key and decoded via a residual reconstruction module, yielding channel-wise predictions with minimal computational overhead.
Objective Optimization and Inference Execution: The training loss integrates reconstruction fidelity, residual sparsity regularization, and statistical alignment objectives. During inference, the same procedure is executed without parameter updates, enabling efficient and accurate multichannel reconstruction.
A.6 Dataset explanations
In experiments, data from different scenarios are used to validate different resource-constrained edge devices, including traffic data (NYC-taxi, DC bike), energy consumption data (Electricity, Solar energy), environmental forecasting data (Weather), and industrial manufacturing monitoring data (Sensor drift). Each dataset contains seven times its periodic length, as indicated in Table 3. Specifically, NYC-taxi (168 timestamps), DC bike (210 timestamps), Electricity (168 timestamps), Solar energy (168 timestamps), Weather (210 timestamps), and Sensor drift (175 timestamps).
Different channels represent unique locations for traffic data (NYC-taxi, DC bike), energy consumption data (Electricity, Solar energy), and environmental forecasting data (Weather), while for industrial manufacturing monitoring data (Sensor drift), the 96 combinations of 16 micro-heated metal oxide (MOX) gas sensors and 6 different target gases are used as channels.
The datasets use a four-times periodic length in each sliding window prediction task, with 3/4 used as the training set and 1/4 used as the testing set. Thus, with a seven-times periodic length, we constructed three sliding windows.
A.7 Model integrations
To demonstrate the compatibility of our compression-decompression framework with different widely used predictors, we integrate the framework with four representative modules: Multilayer Perceptron (MLP), Linear, Transformer, and Convolutional variants. The architectural details of each single-channel predictor are described as follows:
• MLP variant: Built upon the TSMixer-based MLP, this variant incorporates an additional flatten layer, one linear layer, and a ReLU activation, appended to the end of the original structure.
• Transformer variant: Based on the PatchTST encoder, this variant extends the original architecture by adding a flatten layer, two linear layers, and a ReLU activation.
• Linear variant: A linear predictor derived from HDMixer, consisting of multiple linear layers, with a flatten operation inserted before a final fully connected layer.
• Convolution variant: A convolutional module comprising two convolutional layers, followed by a flatten operation, two fully connected layers, and a ReLU activation function.
A.8 Result analysis with MSE
Table 3: MSE comparison across models
@@ A.8 Result analysis with MSE
In terms of MSE, the compression–decompression variants generally achieve performance that is comparable to, or even better than, that of the original models. This can be understood through the Information Bottleneck (IB) framework, which frames compression as a way to retain only task-relevant information while discarding noise. In our setting, the compressed representation preserves the predictive structure of the input, as evidenced by the high Pearson correlation between and the reconstructed prediction. This suggests that the compression process acts as a form of regularization, improving generalization by suppressing irrelevant variability. As a result, the model may achieve a lower MSE than the original.
As shown in Table 3, the best MSE scores for each row are underlined, and the second-best scores are bolded. Our framework’s variants—PCDF-MLP, PCDF-Conv, PCDF-Trans, and PCDF-Linear—frequently produce underlined or bolded values across diverse datasets, demonstrating strong generalization. For instance, PCDF-MLP achieves the best MSE on NYC Taxi (0.16690), and PCDF-Linear records the best or second-best scores on Electricity (0.10462), Sensor Drift (0.22586), and Weather (0.13908). In the DC Bike dataset, PCDF-Trans achieves the best score (0.19908) at 10 channels, while PCDF-Conv obtains the second-best at 20 channels (0.20131). These results demonstrate that our approach remains competitive even in diverse settings with varying channel counts and periodicities.
An exception is observed with TSMixer and its variant. Two potential reasons may explain this phenomenon. First from the perspective of dataset characteristics, performance varies across different data sources. The NYC Taxi and Electricity datasets exhibit clearer seasonal patterns compared to DC Bike, Solar Energy, and Sensor Drift, making them easier to predict and resulting in lower MSE values. Models such as TSMixer perform particularly well on datasets like Electricity. Conversely, in datasets such as Solar Energy, where seasonal patterns are less pronounced, conventional predictors tend to struggle, yielding higher MSE values. In such cases, leveraging the periodic key to extract latent periodic structures leads to improved prediction performance and lower MSE. Second, the information loss introduced by the compression–decompression process may not be fully compensated for by the MLP-based predictor used in this variant.
A.9 Result analysis with runtime
Table 4: Runtime comparison across models. The reported computational time is measured in seconds
@@ A.9 Result analysis with runtime
With respect to computational time, the compression–decompression integrations typically require less runtime than the original models. An exception occurs with the DLinear model, as shown in the appendix. Since DLinear consists of only a single fully connected layer, its computational cost is already extremely low. However, its prediction accuracy is noticeably lower than that of the other models in most cases.
As shown in Table 4, our framework’s variants (PCDF-MLP, PCDF-Conv, PCDF-Trans, and PCDF-Linear) consistently achieve the lowest or second-lowest runtime across all datasets and channel configurations. For example, PCDF-Conv records the lowest time (0.00151s) on Solar Energy (5 channels), and PCDF-Trans leads in speed on NYC Taxi (10 channels, 0.00126s) and DC Bike (20 channels, 0.00234s). Even at higher input dimensions, such as Sensor Drift with 40 channels, PCDF-Linear maintains competitive speed (0.00375s), significantly outperforming attention-based baselines like Autoformer and Transformer-based models by an order of magnitude. These results reinforce that the proposed compression–decompression approach offers not only competitive accuracy but also practical advantages in resource-constrained environments due to its efficiency.

### v2

@@ 6.5. Ablation Analysis
7. Conclusion
This paper proposes a novel predictability-aware compression-decompression framework tailored for multichannel time series prediction. By leveraging a seasonality-guided convolution process with circularly orthogonal key matrices, the framework effectively preserves latent temporal dependencies during compression and mitigates reconstruction errors during decompression. Extensive experiments across diverse datasets and prediction backbones demonstrate that our framework consistently achieves lower runtime and improved prediction accuracy, as evidenced by reduced Cobb–Douglas production indices. Overall, this work offers a flexible and efficient solution for compressing multichannel time series data without sacrificing predictive performance, paving the way for more resource-efficient forecasting in edge–cloud environments.
Limitations and future work.
@@ 7. Conclusion
Acknowledgements.
We acknowledge the support of NSF (2520547). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the NSF.
Appendix A Appendix
A.1. Upper Bound Derivation of the Information Bottleneck Objective
We adopt the Information Bottleneck (IB) principle (Hu et al., 2024), with the objective formulated as:
(14) $\mathcal{L}_{\text{IB}}=I(Y;\ X^{C}_{t:t-L})-\beta I(Y;\ X^{C}_{t:t+H}),$
where $X^{C}_{t:t-L}\in\mathbb{R}^{C\times L}$ is the historical input, $X^{C}_{t:t+H}\in\mathbb{R}^{C\times H}$ is the future target to predict, $Y\in\mathbb{R}^{L}$ is the compressed representation derived from $X^{C}_{t:t-L}$, and $\beta>0$ is a trade-off parameter balancing compression and prediction.
Upper Bound on $I(Y;X^{C}_{t:t-L})$
Following Assumption 2, we assume the historical input includes Gaussian redundancy:
(15) $X^{C}_{t:t-L}=S+\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,\sigma^{2}I_{CL}),$
where $S\in\mathbb{R}^{C\times L}$ is the true signal and $\varepsilon\in\mathbb{R}^{CL}$ is i.i.d. noise. We apply a linear compression:
(16) $Y=KX^{C}_{t:t-L}=KS+K\varepsilon,$
where $K\in\mathbb{R}^{L\times CL}$ is the projection matrix. Assuming a Gaussian encoder, the mutual information upper bound becomes:
(17) $I(Y;\ X^{C}_{t:t-L})\leq\frac{1}{2}\log\left(1+\frac{1}{CL\cdot\sigma^{2}}\right).$
Lower Bound on $I(Y;\ X^{C}_{t:t+H})$
We now bound the prediction term using a variational approximation $q(x\mid y)$ of the true conditional:
$I(Y;\ X^{C}_{t:t+H})$ $=\mathbb{E}_{p(y,\hat{x})}\left[\log q(\hat{x}\mid y)\right]+$
(18) $\underbrace{\mathbb{E}_{p(y,\hat{x})}\left[\log\frac{p(\hat{x}\mid y)}{q(\hat{x}\mid y)}\right]}_{\text{KL divergence}\geq 0}+H(X^{C}_{t:t+H})$
Hence, the variational lower bound is:
(19) $I(Y;\ X^{C}_{t:t+H})\geq\mathbb{E}_{p(y,\hat{x})}\left[\log q(\hat{x}\mid y)\right]+H(X^{C}_{t:t+H}).$
If the model is well trained, $q(\hat{x}\mid y)\approx p(\hat{x}\mid y)$, and:
(20) $I(Y;\ X^{C}_{t:t+H})\approx H(\hat{X}^{C}_{t:t+H})-H(X^{C}_{t:t+H})=0.$
Final Upper Bound
Combining both bounds, we obtain:
(21) $\mathcal{L}_{\text{IB}}\leq\frac{1}{2}\log\left(1+\frac{1}{CL\cdot\sigma^{2}}\right)$
This shows that increasing channel count $C$, history length $L$, or noise variance $\sigma^{2}$ will decrease the IB objective, promoting stronger compression.
A.2. Predictability Preservation in Compression
We provide a formal justification for how circular convolution preserves temporal predictability when the key structure aligns with the latent seasonality of the input.
Let each channel $X^{i}\in\mathbb{R}^{L}$ be partitioned into $m$ non-overlapping segments of length $\tau$, such that $L=m\tau$. Let the key $K\in\mathbb{R}^{\tau}$ be a seasonal vector with period $\tau$, extended to length $L$ via circular repetition.
Each compressed channel is encoded as:
(22) $Y^{i}[t]$ $=\text{Encode}(K,X^{i})[t]=\sum_{z=0}^{L-1}x^{i}_{z}\cdot k_{(t-z)\bmod L},$
$t\in\{0,1,\dots,L-1\}.$
We can express the first $\tau$ outputs compactly as:
(23) $Y^{i}_{0:\tau}=\begin{bmatrix}K^{\langle 0\rangle}\odot X^{i}\\ K^{\langle 1\rangle}\odot X^{i}\\ \vdots\\ K^{\langle\tau-1\rangle}\odot X^{i}\end{bmatrix}$,
where $K^{\langle t\rangle}$ denotes the circular shift of $K$ by $t$ positions, and $\odot$ denotes element-wise multiplication followed by summation.
To assess the preservation of seasonal structure, we compute the Pearson correlation coefficient between $Y^{i}[t]$ and its delayed version $Y^{i}[t+\tau]$:
(24) $R=$
$\frac{\sum_{t=0}^{\tau-1}\left(Y^{i}[t]-\bar{Y}^{i}_{0:\tau}\right)\left(Y^{i}[t+\tau]-\bar{Y}^{i}_{\tau:2\tau}\right)}{\sqrt{\sum_{t=0}^{\tau-1}\left(Y^{i}[t]-\bar{Y}^{i}_{0:\tau}\right)^{2}}\cdot\sqrt{\sum_{t=0}^{\tau-1}\left(Y^{i}[t+\tau]-\bar{Y}^{i}_{\tau:2\tau}\right)^{2}}},$
where $\bar{Y}^{i}_{0:\tau}$ and $\bar{Y}^{i}_{\tau:2\tau}$ denote the mean values of the first and second seasonal segments of the encoded signal $Y^{i}$, respectively. This formulation evaluates the seasonal correlation between two adjacent cycles in the compressed output. Then $K^{\langle t\rangle}$ = $K^{\langle t+\tau\rangle}$:
(25) $R=\frac{\sum_{t=0}^{\tau-1}(Y^{i}[t]-\bar{Y}^{i})^{2}}{\sqrt{\sum(Y^{i}[t]-\bar{Y}^{i})^{2}}\cdot\sqrt{\sum(Y^{i}[t]-\bar{Y}^{i})^{2}}}=1.$
Hence, when the circular key aligns with the latent seasonal period $\tau$, the circular convolution operation effectively preserves temporal dependencies in the encoded output, ensuring the retention of predictive information.
Table 5. MSE comparison across models
@@ A.2. Predictability Preservation in Compression
Table 6. Runtime comparison across models. The reported computational time is measured in seconds
@@
A.3. Time Complexity Justification
(26) $\mathcal{O}(L)_{\text{comp}}$ $=\underbrace{DEL^{2}}_{\text{Prediction}}+\underbrace{\frac{CL^{2}}{\tau}+2L\log L}_{\text{Compression}}$
$\quad+\underbrace{L+3CL+CL^{2}+CL^{2}}_{\text{Decompression}}$
$\approx DEL^{2}+\frac{CL^{2}}{\tau}+CL^{2},$
where $E$ denotes the number of training epochs, $L$ is the number of timestamps, $C$ $(C<L)$ is the number of channels, $D$ is the depth of the predictor, and $\tau$ is the least common multiple (LCM) seasonal period across different channels. In the asymptotic approximation, we neglect lower-order terms.
The time complexity of the decompression module, dominated by residual blocks, can be broken down into standard neural network operations:
(27) $\mathcal{O}(L)_{\text{decompression}}$ $=L+3CL+CL^{2}+CL^{2}$
In contrast, a naive multichannel predictor that processes all channels jointly without compression incurs the following computational complexity:
(28) $\mathcal{O}(L)_{\text{naive}}=DECL^{2}$
To determine the minimum number of channels $C$ for which the proposed method achieves lower computational cost than the naive baseline, we compare both complexities:
(29) $\mathcal{O}(L)_{\text{naive}}$ $>\mathcal{O}(L)_{\text{comp}}$
$DECL^{2}$ $>DEL^{2}+CL^{2}+\frac{CL^{2}}{\tau}$
$\Rightarrow\quad C$ $>\frac{DE}{DE-1-\frac{1}{\tau}}$
Under the assumptions that $DE\gg 1$ and $\frac{1}{\tau}\to 0$, the lower bound simplifies to:
(30) $C>1$
This result indicates that the proposed compression-based framework offers a computational advantage when the number of channels exceeds one. Furthermore, since the additional costs of compression and decompression are subdominant in the asymptotic regime, they do not alter the scaling benefit of the proposed method.
A.4. Framework Pipeline with Pseudocode
Pseudocode for training process.
The pseudocode in Algorithm 1 describes the training phase of our compression–prediction-decompression framework.
Algorithm 1 Training Phase of Compression–Prediction Framework
1:for epoch =1=1 to $T$ do
2:  Compression:
3:  for $i=1$ to $C$ do
4:   Split $X^{i}\in\mathbb{R}^{L}$ into segments of length $\tau$: $X^{i}_{0:\tau},X^{i}_{\tau:2\tau},\dots$
5:   for each complete segment do
6:    $Y^{i}\leftarrow\text{Encode}(X^{i}_{j\tau:(j+1)\tau},K)$
7:   end for
8:  end for
9:  $Y\leftarrow\sum_{i=1}^{C}Y^{i}$
10:  Prediction:
11:  $\hat{Y}\leftarrow\text{Predict}(Y)$
12:  Decompression:
[…]
14:  $\hat{X}\leftarrow\text{Dense}\left(\frac{\text{Copy}(\tilde{X})}{\sum k^{2}}+\text{Residual}\left(\frac{\text{Copy}(\tilde{X})}{\sum k^{2}}\right)\right)$
15:  Loss Computation:
16:  Compute $\mathcal{L}$ as defined in Eq. 31
17:  Parameter Update:
18:  $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}$
19:end for
Additional techniques for training the model
Following Dish-TS (Fan and others, 2023), we apply normalization after compression and denormalization before decompression.
Training loss functions
The total training objective consists of three terms:
(31) $\mathcal{L}=$ $\underbrace{\|\hat{X}-X\|^{2}}_{\text{Prediction Loss}}+\alpha\underbrace{\left(\sum_{i=1}^{C}|\text{Residual}_{i}|-\sum_{i=1}^{C}\hat{X}_{i}\right)^{2}}_{\text{Regulation Term}}$
$+\beta\underbrace{\|\varphi_{Y}-\varphi_{\hat{Y}}\|^{2}}_{\text{Latent Consistency}}.$
Here, $\text{Residual}_{i}$ represents the residual correction output for the $i$-th channel, $\varphi_{Y}$ and $\varphi_{\hat{Y}}$ denote the mean values of the original and predicted compressed representations, respectively. This latent consistency term encourages the predictor to preserve global statistics (e.g., mean trends) in the compressed space, ensuring more coherent reconstructions after decompression. The regulation term encourages sparsity and suppresses overfitting.
To improve numerical stability, we use scale-invariant gradient clipping (Li and others, 2022). For each parameter $w$ and its gradient $g$, we clip the gradient as:
(32) $g\leftarrow g\cdot\min\left(1,\ \frac{\alpha\|w\|}{\|g\|}\right),$
where $\alpha>0$ is a hyperparameter, $g$ denotes the gradient and $w$ is the corresponding model parameter.
A.5. Result Analysis with MSE
As shown in Table 5, the best scores are underlined and the second-best are bolded. The PCDF variants (MLP, Conv, Trans, and Linear) frequently achieve top results across datasets.
A.6. Result Analysis with Runtime
As shown in Table 6, our variants (PCDF-MLP, PCDF-Conv, PCDF-Trans, PCDF-Linear) consistently achieve the lowest or second-lowest runtime across datasets and channel settings. These results highlight both the accuracy and efficiency of our approach, especially in resource-limited settings.